4. (A) (25 marks) The Chamber of Commerce is planning its July 4 gala ball. There are two a. Country Golf Club, which has a fixed rental cost of $2.000 plus a charge of $80 per b. Town Hall, which has a fixed rental cost of $6,600 The Chamber of Commerce can possible venues: person for its own catering of meals and serving of drinks. hire a caterer for meals and waiters to serve drinks at $60 per person. Compute breakeven point for each venue in terms of tickets sold. (B) (25 marks) Suppose our corporations breakeven point is dales of 1,000,000 TL and our fixed costs are 400,000 TL a. b. Compute contribution margin percentage Compute the selling price if variable costs are 12 TL per unit Part b
Answer #1

4 (A) Let the break even point (in terms of no. tickets) be "x"

Thus total cost for Country Golf Club = 2000+80x and total cost for Town Hall = 6600+60x

At break even both the costs will be the same. Thus 2000+80x = 6600+60x

or 4600=20x

0r x = 230

Thus breakeven point = 230 tickets

(B): a: Breakeven = fixed costs/(selling price per unit - variable costs per unit)

So, 1,000,000 TL = 400,000 TL/(selling price per unit - variable costs per unit)

or selling price per unit - variable cost per unit = 400,000/1,000,000 = 0.4

Or contribution margin% = 0.4*100 = 40%

b: Now (sales - VC)/sales = 0.4

or (sales-12)/sales = 0.4

or sales - 12 = 0.4 sales

or 0.6 sales = 12

or sales = 12/0.6 = 20

Thus selling price = 20 TL per unit